OpenStudy (beccasmith):

WILL MEDAL IF CORRECT!!! Morgan lives 4/5 of a mile from the park. Stanley lives 1/3 of that distance. How far does Stanley live from the park? (P.S. Please answer in the simplest form possible.)

OpenStudy (anonymous):

do you have to get a common denominator?

OpenStudy (bibby):

in general, of means multiplication in math

OpenStudy (bibby):

\(\large\frac{1}{3}of\frac{4}{5}=\frac{1}{3}*\frac{4}{5}=\)

OpenStudy (beccasmith):

Is 4/15 the correct answer?

OpenStudy (bibby):

probably

OpenStudy (anonymous):

thats what i got

OpenStudy (beccasmith):

So it's the simplest form? Sorry I am terrible at math.

OpenStudy (bibby):

you're both right, yeah

OpenStudy (bibby):

can you simplify it further? do 4 and 15 have a common factor

OpenStudy (beccasmith):

Thank you.I believe not.

OpenStudy (beccasmith):

Who would like the medal?

OpenStudy (beccasmith):

I'll give it to bibby

OpenStudy (bibby):

I've got like 10 medals she can have it

OpenStudy (bibby):

1600 medals* I'll survive